### v1

# Many-Body Rashba Spin-Orbit Interaction and Exciton Spin Relaxation in Atomically Thin Semiconductor Structures

Abstract
We propose a previously unexplored spin-orbit interaction mechanism by establishing a mesoscopic many-particle Rashba Hamiltonian. In lowest order, this Hamiltonian self-consistently describes exciton spin relaxation in monolayer transition metal dichalcogenides (TMDC) due to local electric fields caused by spatial asymmetries in the dielectric environment. For a monolayer MoSe2 on a SiO2 substrate above 77 K showing a meV bright-dark splitting, the local electric field causes fast intravalley spin relaxation on a sub-picosecond timescale, whereas it is negligible for other TMDCs with larger bright-dark splitting.
preprint: APS/123-QED
I Introduction
Atomically thin semiconductors Wang et al. (2018a) host a broad variety of incoherent exciton scattering mechanisms. Based on the valley-selective optical selection rules, next to spin-conserving exciton-phonon scattering Selig et al. (2018); Brem et al. (2020); Wallauer et al. (2021); Katzer et al. (2023), Coulomb scattering mechanisms such as spin-conserving Dexter interaction Berghäuser et al. (2018) and spin-nonconserving intra- and intervalley exchange Guo et al. (2019); Maialle et al. (1993); Yu and Wu (2014); Selig et al. (2019, 2020); Kwong et al. (2021) via double spin flips are relevant. Moreover, single spin flips can occur via, e.g., Rashba interaction. Rashba spin-orbit interaction can already affect linear optical spectra of a spin-forbidden mixture of $s$-$p$ orbital excitons in monolayers Cao et al. (2024, 2025). However, these signatures are very weak and have not been measured until now. For $s$-orbital excitonic states, which are usually measured in absorption or photoluminescence due to their large optical oscillator strength, no brightening of spin-dark excitonic states in out-of-plane electric fields can occur, since Rashba spin coupling vanishes at zero center-of-mass (COM) momentum within the light cone Mittenzwey et al. (2025). Thus, Rashba spin coupling is only relevant for incoherent $s$-orbital excitons occupying COM momenta well beyond the light cone. In contrast to Coulomb-exchange-induced double spin flips, single spin flip scattering mechanisms are less understood, as they are often included via external parameters Ochoa and Roldán (2013); Dery and Song (2015); Song and Dery (2013); Yang et al. (2020) or resulting from a complex band structure Wang et al. (2018b); Molina-Sánchez et al. (2017). Recently, chiral phonons have been identified as an important single spin-flip mechanism Lagarde et al. (2024); Chan et al. (2025); Zhang and Cao (2022).
In this work, we derive a new many-body Rashba spin-orbit interaction mechanism Winkler (2003); Bychkov and Rashba (1984); Kormányos et al. (2014):
$H_{\text{BR}}=-\frac{\alpha_{\text{BR}}}{\hbar}\bm{\sigma}\cdot\left(\mathbf{E}(\mathbf{r},z)\times\mathbf{p}\right),$ (1)
which results from self-consistently described out-of-plane electric fields $\mathbf{E}(\mathbf{r},z)$, induced by the dielectric environment. Here, $\bm{\sigma}=\begin{pmatrix}\sigma_{x}&\sigma_{y}&\sigma_{z}\end{pmatrix}^{\top}$ is the vector of the Pauli matrices $\sigma_{i}$, $\mathbf{p}=\frac{\hbar}{\mathrm{i}}\nabla$ is the momentum operator and $\alpha_{\text{BR}}$ is the Rashba coupling constant, which, in solids, is greatly enhanced compared to the atomic constant appearing in the Pauli equation Kormányos et al. (2014); Winkler (2003).
Starting from Eq. (1), we derive a second-quantized description of the Rashba interaction induced by the optically induced (image) charge density due to spatial asymmetries in the dielectric environment of an atomically thin semiconductor (Sec. II), which is subsequently transformed into an excitonic description (Sec. III). In combination with exciton-phonon interaction generating non-zero COM-momentum excitons (Sec. IV), the described spin-orbit interaction leads to a phonon-assisted exciton spin relaxation. We illustrate the new mechanism for a monolayer MoSe2 and quantify its impact on the incoherent exciton dynamics via microscopic calculations (Sec. V).
II Electron-Hole Many-Particle Rashba Hamiltonian due to Self-Consistent Fields in an Inhomogeneous Dielectric Environment
We consider the following geometry, cf. Fig. 1: A monolayer TMDC ($\epsilon_{s}$) is sandwiched between a substrate ($\epsilon_{1}$) and superstrate ($\epsilon_{2}$) dielectric material with a small vacuum gap in between due to van der Waals interaction Florian et al. (2018); Drüppel et al. (2017). An induced charge density $\hat{\rho}_{\mathbf{q}}$ via, e.g., optical excitation, creates image charges, which induce a net out-of-plane electric field $\hat{E}_{z,\mathbf{q}}$ in an inhomogeneous dielectric environment ($\epsilon_{1}\neq\epsilon_{2}$).
To investigate the consequences thereof, we employ the method of second quantization for Bloch electrons Malic and Knorr (2013). The Rashba Hamiltonian from Eq. (1) reads:
$\hat{H}_{\text{BR}}=$ $\,\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{q},\\ \lambda_{1},\lambda_{2},\\ \xi,\xi^{\prime},s\end{subarray}}\hat{E}_{z,\mathbf{q}+\mathbf{K}^{\xi^{\prime}}-\mathbf{K}^{\xi}}S_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda_{1},\lambda_{2},\xi^{\prime},\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{1},\mathbf{k}+\mathbf{q}}^{\xi^{\prime},\bar{s}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{2},\mathbf{k}}^{\xi,s}.$ (2)
A detailed derivation of Eq. 2 is given in App. A. Here, $\hat{E}_{z,\mathbf{q}}$ is the spatially inhomogeneous, quantum-confined out-of-plane ($z$-direction) oriented electric field, cf. Fig. 1, defined as:
$\hat{E}_{z,\mathbf{q}}=\frac{1}{\mathcal{A}}\int\mathrm{d}z\,\left|\xi(z)\right|^{2}\hat{E}_{z,\mathbf{q}}(z),$ (3)
[…]
$\Psi_{\mathbf{k},\lambda}^{\xi,s}(\mathbf{r},z)=\frac{1}{\sqrt{\mathcal{A}}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}u_{\lambda,\mathbf{k}}^{\xi,s}(\mathbf{r})\xi(z)\chi_{s},$ (4)
with Bloch factor $u_{\lambda,\mathbf{k}}^{\xi,s}(\mathbf{r})$ and spin wave function $\chi_{s}$ and is defined as:
[…]
where $\mathbf{p}_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda,\lambda^{\prime},\xi^{\prime},\xi,\bar{s},s}$ is the momentum matrix element, cf. Eq. (34) in App. A. At first glance, Eq. (LABEL:eq:RashbaMatrixElementGeneral), and therefore Eq. (2), seem to break the rotational symmetry in the $x$-$y$-plane. Actually, this is not the case, as a rotation in the $x$-$y$-plane simultaneously acts on both the spin vector $\bm{\sigma}$ and the cross product of field and momentum $\mathbf{E}(\mathbf{r},z)\times\mathbf{p}$, which always ensures rotational invariance of the Rashba Hamiltonian in Eq. (2). In Eq. (2), an external electric field $\hat{E}_{z,\mathbf{q}}(z)$ can be applied in $z$-direction with vanishing in-plane variation, so that $\hat{E}_{z,\mathbf{q}}(z)\equiv E_{z}^{\text{ext}}\mathcal{A}\delta_{\mathbf{q},\mathbf{0}}$, where $E_{z}^{\text{ext}}$ can be generated by, e.g., static electrical gating Kovalchuk et al. (2023), or it can be an in-plane propagating optical field Wang et al. (2017) with an almost vanishing in-plane photon momentum. Static fields couple mainly via the intraband Rashba interaction with $\lambda_{1}=\lambda_{2}$, whereas optical fields couple mainly via the interband Rashba interaction with $\lambda_{1}\neq\lambda_{2}$.
Figure 1: Sketch of the sample geometry. $\epsilon_{1/2}$ is the static dielectric constant of substrate/superstrate and $\epsilon_{s}$ is the dielectric constant of the thin semiconductor, which are separated by a vacuum gap of width $h$. $\hat{\rho}_{\mathbf{q}}$, cf. Eq. (7), is the charge density and $\hat{E}_{z,\mathbf{q}}$ is the (operator-valued) electric field induced by image charges at the boundaries, cf. Eq. (LABEL:eq:ElectricFieldConfined).
In the case considered here, i.e., without electrical gating or in-plane propagating optical fields, the electric field is induced by a self-arranging charge-density fluctuation of electrons and holes $\hat{\rho}_{\mathbf{q}}(z)$ within the atomically thin semiconductor itself. The occurring charge distribution $\hat{\rho}_{\mathbf{q}}(z)$ results from the dielectric boundary conditions for the longitudinal fields $\mathbf{E}^{\parallel}$ and $\mathbf{D}^{\parallel}$ implied by the macroscopic Maxwell equations. The boundary conditions for the normal vector $\mathbf{n}$ in $z$-direction read: $\mathbf{n}\times\mathbf{E}_{i}^{\parallel}=\mathbf{n}\times\mathbf{E}_{j}^{\parallel}$ and $\mathbf{n}\cdot\mathbf{D}_{i}^{\parallel}=\mathbf{n}\cdot\mathbf{D}_{j}^{\parallel}$ at the interface, cf. Fig. 1, between layer $i$ and layer $j$ inducing electronic charges $\hat{\rho}_{\mathbf{q}}$ in the heterostructure, that have to be adjusted self-consistently. To account for this self-consistency in the charge $\hat{\rho}_{\mathbf{q}}$, we solve the inhomogeneous Poisson equation for a layered dielectric environment with the operator-valued charge density $\hat{\rho}_{\mathbf{q}}(z)$ as a source. The solution can be obtained by the Green’s function $G_{\mathbf{q}}(z,z^{\prime})$ of Poisson’s equation. The local, quantum-confined electric field can be expressed as, cf. App. B:
$\begin{split}&\hat{E}_{z,\mathbf{q}}=-\frac{1}{\mathcal{A}}\int\mathrm{d}z\,\left|\xi(z)\right|^{2}\partial_{z}\int\mathrm{d}z^{\prime}G_{\mathbf{q}}(z,z^{\prime})\hat{\rho}_{\mathbf{q}}(z^{\prime}).\end{split}$ (6)
The occurring source charge density reads:
$\hat{\rho}_{\mathbf{q}}(z)=$ $\,-e\left|\xi(z)\right|^{2}\sum_{\begin{subarray}{c}\mathbf{k},\xi,\xi^{\prime},s,\\ \lambda,\lambda^{\prime}\end{subarray}}F_{\lambda,\lambda^{\prime},\mathbf{k}-\mathbf{q},\mathbf{k}}^{\xi,\xi^{\prime},s,s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda,\mathbf{k}-\mathbf{q}}^{\xi,s}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda^{\prime},\mathbf{k}}^{\xi^{\prime},s},$ (7)
where $e$ is the elementary charge and $F_{\lambda,\lambda^{\prime},\mathbf{k}-\mathbf{q},\mathbf{k}}^{\xi,\xi^{\prime},s,s}$ is the form factor, given by:
$F_{\lambda,\lambda^{\prime},\mathbf{k}-\mathbf{q},\mathbf{k}}^{\xi,\xi^{\prime},s,s}=\frac{1}{\mathcal{A}_{\text{UC}}}\int_{\mathcal{A}_{\text{UC}}}\mathrm{d}^{2}r\,\,u^{*}\vphantom{u}_{\lambda,\mathbf{k}-\mathbf{q}}{}^{\xi,s}(\mathbf{r})u_{\lambda^{\prime},\mathbf{k}}^{\xi^{\prime},s}(\mathbf{r}),$ (8)
with unit cell area $\mathcal{A}_{\text{UC}}$ and integration over the lattice-periodic Bloch factor $u_{\lambda,\mathbf{k}}^{\xi,s}(\mathbf{r})$. In contrast to Ref. Mittenzwey et al. (2025), where the charge density is treated in a classical mean-field limit, the crucial point of the present work is to preserve the operator-valued character and thus quantum fluctuations in Eq. (7). This way, not only Hartree contributions but also Fock- or higher correlations can be considered and improve the treatment of many-body interactions. Using Eqs. (LABEL:eq:ElectricFieldConfined)–(8), we arrive at the following form of the Rashba Hamiltonian from Eq. (2):
$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\\ \lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\\ \xi,\xi^{\prime},\xi^{\prime\prime},s,s^{\prime}\end{subarray}}E_{z,\mathbf{q}+\mathbf{K}^{\xi^{\prime}}-\mathbf{K}^{\xi}}^{\text{loc}}F_{\lambda_{2},\lambda_{3},\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},\xi^{\prime\prime},s^{\prime},s^{\prime}}\\ &\,\quad\times S_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda_{1},\lambda_{4},\xi^{\prime},\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{1},\mathbf{k}+\mathbf{q}}^{\xi^{\prime},\bar{s}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{2},\mathbf{k}^{\prime}-\mathbf{q}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{3},\mathbf{k}^{\prime}}^{\xi^{\prime\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{4},\mathbf{k}}^{\xi,s}.\end{split}$ (9)
Eq. (LABEL:eq:RashbaHamiltonianManyBody) is the full many-body Rashba Hamiltonian, which, at this level, accounts for various single spin flip scattering $s\rightarrow\bar{s}$ analog to the many-body Coulomb Hamiltonian, e.g. intraband ($\lambda_{1}=\lambda_{4}$ and $\lambda_{2}=\lambda_{3}$), interband ($\lambda_{1}\neq\lambda_{4}$ and $\lambda_{2}\neq\lambda_{3}$) and Auger-like ($\lambda_{1}=\lambda_{4}$ and $\lambda_{2}\neq\lambda_{3}$ or vice versa) electron-electron, hole-hole and electron-hole intravalley ($\xi=\xi^{\prime}$) and intervalley ($\xi^{\prime}\neq\xi$) scattering. As mediating quantity, in Eq. (LABEL:eq:RashbaHamiltonianManyBody), the (classical) local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ can be identified:
$\begin{split}E_{z,\mathbf{q}}^{\text{loc}}=\frac{e}{\mathcal{A}2\epsilon_{0}\epsilon_{s,\bot}}\frac{g_{\mathbf{q}}}{f_{\mathbf{q}}},\end{split}$ (10)
where $\epsilon_{0}$ is the vacuum permittivity and $\epsilon_{s,\bot}$ is the relative out-of-plane dielectric constant of the thin semiconductor. The unitless functions $g_{\mathbf{q}}$ and $f_{\mathbf{q}}$, which encode the momentum dependence determined by the dielectric structure, are given in Eq. (41) and Eq. (42) in App. B, respectively. The electric field in Eq. (10) stems solely from the homogeneous solution of the Poisson equation, which arises due to image charges determined by the electrostatic boundary conditions. To evaluate Eq. (LABEL:eq:RashbaHamiltonianManyBody), we assume a perfectly flat TMDC monolayer with a $z$-dependent dielectric environment $\epsilon(z)$ without any vacancies Latini et al. (2015). Consequently, if we set $\epsilon_{1}=\epsilon_{2}$, i.e., equal substrate and superstrate material in Fig. 1, the field in Eq. (10) vanishes:
$E_{z,\mathbf{q}}^{\text{loc}}\Big{|}_{\epsilon_{1}=\epsilon_{2}}=0.$ (11)
We note, that this is not always the case, since, e.g., chalcogen vacancies can induce spatial asymmetries in $z$-direction in the charge distribution Özcan et al. (2024); Min et al. (2024).
On the other hand, if substrate and superstrate differ in their respective dielectric constants $\epsilon_{1}\neq\epsilon_{2}$, the local electric field is nonzero:
$E_{z,\mathbf{q}}^{\text{loc}}\Big{|}_{\epsilon_{1}\neq\epsilon_{2}}\neq 0.$ (12)
This is the situation depicted in Fig. 1.
Figure 2: Local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ from Eq. (10) with normalization area $\mathcal{A}=1\,\text{nm}^{2}$ for an example MoSe2 monolayer ($\epsilon_{s,\perp}=7.2$ Laturia et al. (2018)) on a SiO2 substrate ($\epsilon_{1}=3.9$ Xue et al. (2011), $\epsilon_{2}=1$) and on a sapphire substrate ($\epsilon_{1}=\sqrt{11.6\cdot 9.4}$ Fontanella et al. (1974), $\epsilon_{2}=1$) for a substrate distance $h$ of 0 nm and 0.3 nm.
In Fig. 2, we depict the local electric field in Eq. (10) for an example MoSe2 monolayer on a SiO2 and sapphire substrate for different substrate distances. Here, the local electric field is well-localized in momentum space and reaches values larger than 1 V nm-1 at $\mathbf{q}=\mathbf{0}$ on a sapphire substrate. At first glance, the calculated fields are in the range of V nm-1 and appear to be rather strong. However, the strength of the local electric field cannot be directly compared to the total field acting in the sample, since it acts under a momentum sum in Eq. (LABEL:eq:RashbaHamiltonianManyBody), so that the overall impact of the local electric field is provided by the momentum integral of the curves in Fig. 2 times the semiconductor area $\mathcal{A}$. Moreover, a distance $h$ between the atomically thin semiconductor and the substrate, which always occurs to some degree in TMDC samples on substrates on the order of h=0.3h=0.3\,nm–0.50.5\,nm Florian et al. (2018); Drüppel et al. (2017), affects the localization. The larger the distance $h$, the stronger the electric field is localized around zero momentum resulting in a decreased overall impact of the local electric field on the electrons and holes. Due to the relatively strong localization around small in-plane momenta $\mathbf{q}$, intervalley scattering occurring due to the derived Rashba interaction, cf. Eq. (LABEL:eq:RashbaHamiltonianManyBody), at large valley momenta $|\mathbf{q}|=K=\frac{4\pi}{3a_{0}}=12.6\,$nm-1 ($K$ valley-momentum transfer with lattice constant $a_{0}$ Kormányos et al. (2015)) is strongly suppressed.
III Excitonic Rashba Hamiltonian
To illustrate the basic impact of the spin-flip processes introduced by the charge- and field fluctuations of the dielectric environment, we restrict the analysis to intraband electron or hole scattering mediated by the Rashba Hamiltonian in Eq. (LABEL:eq:RashbaHamiltonianManyBody). This process is related to the so-called direct Coulomb interaction, which constitutes a fundamental scattering process in semiconductors Rohlfing and Louie (2000). For optical excitation at the $K$ points, we apply the low-wavenumber approximation to the scattering matrix elements in Eq. (LABEL:eq:RashbaHamiltonianManyBody), which is valid for small momentum transfers $\mathbf{q}$, and obtain:
$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\lambda,\lambda^{\prime}\\ \xi,\xi^{\prime},s,s^{\prime}\end{subarray}}E_{z,\mathbf{q}}^{\text{loc}}S_{\mathbf{k}}^{\lambda,\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda,\mathbf{k}+\mathbf{q}}^{\xi,\bar{s}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda^{\prime},\mathbf{k}^{\prime}-\mathbf{q}}^{\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda^{\prime},\mathbf{k}^{\prime}}^{\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda,\mathbf{k}}^{\xi,s},\end{split}$ (13)
where $S_{\mathbf{k}}^{\lambda,\xi,\bar{s},s}\equiv S_{\mathbf{k},\mathbf{k}}^{\lambda,\lambda,\xi,\xi,\bar{s},s}$ from Eq. (LABEL:eq:RashbaMatrixElementGeneral). In the excitonic picture, using the approach and definitions in Ref. Katsch et al. (2018), the Hamiltonian in Eq. (LABEL:eq:RashbaHamiltonianManyBodySmallQ) transforms as follows:
$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mu,\nu,\mathbf{Q},\\ \xi,\xi^{\prime},s,s^{\prime}\end{subarray}}\left(S_{\mu,\nu,\mathbf{Q}}^{h,\xi,\xi^{\prime},\bar{s},s^{\prime}}{{}\hat{P}^{\dagger}}{\vphantom{P}}_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}{{}\hat{P}^{\vphantom{\dagger}}}{\vphantom{P}}_{\nu,\mathbf{Q}}^{\xi,\xi^{\prime},\bar{s},s^{\prime}}\right.\\ &\,\left.\quad\quad\quad\quad\quad-S_{\mu,\nu,\mathbf{Q}}^{e,\xi,\xi^{\prime},s,\bar{s}^{\prime}}{{}\hat{P}^{\dagger}}{\vphantom{P}}_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}{{}\hat{P}^{\vphantom{\dagger}}}{\vphantom{P}}_{\nu,\mathbf{Q}}^{\xi,\xi^{\prime},s,\bar{s}^{\prime}}\right),\end{split}$ (14)
which closely resembles the Hamiltonian describing the Rashba interaction on interlayer excitons derived in Ref. Mittenzwey et al. (2025). In contrast to our many-body description in this work, however, the Rashba interaction in Ref. Mittenzwey et al. (2025) originates from a Hartree-limit. In Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton), $\hat{P}^{(\dagger)}\vphantom{P}_{\mu,\mathbf{Q}}{}^{\xi,\xi^{\prime},s,s^{\prime}}$ are the excitonic annihilation (creation) operators, $\mu$ is the excitonic quantum number and $\mathbf{Q}$ is the center-of-mass momentum. Regarding the index pairs of valley $\xi,\xi^{\prime}$ and spin $s,s^{\prime}$, the first index always belongs to the hole and the second to the electron.
The many-particle Rashba Hamiltonian in Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton) encodes single hole spin-flip processes between spin-bright A and spin-dark B excitons, cf. Fig. 3 and between spin-dark A and spin-bright B excitons, cf. Fig. 3, as well as single electron spin flip processes between spin-bright A and spin-dark A excitons, cf. Fig. 3, and between spin-bright B and spin-dark B excitons, cf. Fig. 3. Note, that, in Fig. 3, we only depict the Rashba intravalley spin-flip processes for intravalley excitons at the $K$ valley with $\xi=\xi^{\prime}=K$, while intravalley spin-flip processes are equally possible for intravalley excitons at the $K^{\prime}$ valley with $\xi=\xi^{\prime}=K^{\prime}$ or for any intervalley configuration with $\xi\neq\xi^{\prime}$, e.g., for $\xi=K$ and $\xi^{\prime}=K^{\prime}$.
Figure 3: Scheme of hole spin coupling between spin-bright A and spin-dark B excitons (a), hole spin coupling between spin-dark A and spin-bright B excitons (b), electron spin coupling between spin-bright and spin-dark A excitons (c) and electron spin coupling between spin-dark B and spin-bright B excitons (d) at the $K$ valley induced by the corresponding Rashba coupling matrix elements $S_{\mathbf{Q}}^{h/e,K,K}$ in Eq. (15) and Eq. (16).
The corresponding excitonic Rashba matrix elements occurring in Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton) for single hole spin flips, cf. Fig. 3 and Fig. 3, read:
$\begin{split}S_{\mu,\nu,\mathbf{Q}}^{h,\xi,\xi^{\prime},\bar{s},s^{\prime}}=\alpha_{\text{BR},v}^{\xi}\sum_{\mathbf{q},\mathbf{q}^{\prime}}E_{z,\mathbf{q}-\mathbf{q}^{\prime}}^{h,\text{tot}}\left(\mathrm{i}\left(\delta_{s,\uparrow}-\delta_{s,\downarrow}\right)q_{x}-q_{y}\right)\\ \times\varphi^{*}{\vphantom{\varphi}}_{\mu,\mathbf{q}+\beta_{\xi,\xi^{\prime}}^{s,s^{\prime}}\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}\varphi_{\nu,\mathbf{q}^{\prime}+\beta_{\xi,\xi^{\prime}}^{s,s^{\prime}}\mathbf{Q}}^{\xi,\xi^{\prime},\bar{s},s^{\prime}},\end{split}$ (15)
the excitonic Rashba matrix elements for single electron spin flips, cf. Fig. 3 and Fig. 3, read:
$\begin{split}S_{\mu,\nu,\mathbf{Q}}^{e,\xi,\xi^{\prime},s,\bar{s}^{\prime}}=\alpha_{\text{BR},c}^{\xi^{\prime}}\sum_{\mathbf{q},\mathbf{q}^{\prime}}E_{z,\mathbf{q}-\mathbf{q}^{\prime}}^{e,\text{tot}}\left(\mathrm{i}\left(\delta_{\bar{s}^{\prime},\uparrow}-\delta_{\bar{s}^{\prime},\downarrow}\right)q_{x}-q_{y}\right)\\ \times\varphi^{*}{\vphantom{\varphi}}_{\mu,\mathbf{q}^{\prime}-\alpha_{\xi,\xi^{\prime}}^{s,s^{\prime}}\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}\varphi_{\nu,\mathbf{q}-\alpha_{\xi,\xi^{\prime}}^{s,s^{\prime}}\mathbf{Q}}^{\xi,\xi^{\prime},s,\bar{s}^{\prime}}.\end{split}$ (16)
Here, $\alpha_{\text{BR},v/c}^{\xi}$ is the Rashba coupling constant for holes/electrons at high-symmetry point $\xi$ Kormányos et al. (2014), $\varphi_{\nu,\mathbf{q}}^{\xi,\xi^{\prime},s,s^{\prime}}$ is the excitonic wave function as a solution of the Wannier equation, $\mathbf{q}$ is the relative momentum and $\alpha$, $\beta$ are the effective-mass ratios of the corresponding exciton configuration Katsch et al. (2018). We rewrote Eq. (15) and Eq. (16) in a way that they can address a more general situation compared to the dielectric environment only: $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ can account for the total electric field having internal (local) contributions due to the dielectric environment and external contributions from, e.g., electrical gating, cf. Fig. 4. This way, the excitonic Rashba Hamiltonian in Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton) combines two physically distinct fields, a local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ due to many-particle interactions and an arbitrary in-plane spatially homogeneous external electric field $E_{z}^{\text{ext}}$, in one consistent description, which is described by a total electric field $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ entering Eq. (15) and Eq. (16):
$E_{z,\mathbf{q}}^{h/e,\text{tot}}=E_{z}^{\text{ext}}\delta_{\mathbf{q},\mathbf{0}}\,{+}{/}{-}\,E_{z,\mathbf{q}}^{\text{loc}}.$ (17)
The external electric field $E_{z}^{\text{ext}}$ and the local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ acting on the corresponding electron/hole may amplify or counteract each other, depending on the dielectric environment and the direction of $E_{z}^{\text{ext}}$ .
Figure 4: Alignment of the respective local electric field $E_{z,\mathbf{q}}$ acting on the individual hole and electron of an exciton (shaded ellipse) in a thin semiconductor $\epsilon_{s}$ for different dielectric environments $\epsilon_{1}$ and $\epsilon_{2}$. $E_{z}^{\text{ext}}$ is an external electric field chosen as polarized in positive $z$-direction acting on both electrons and holes.
To examine this aspect in more detail, in Fig. 4, we depict the alignment of the local electric field acting on the corresponding hole and electron of the exciton resulting from Eq. (17) in comparison to an arbitrary external electric field polarized in positive $z$-direction. Here, we observe two major features: The local electric fields acting on electrons and holes entail – beyond their attraction (blue ellipse) – dielectric fields which depend on differences in the dielectric environment $\epsilon_{1}$ and $\epsilon_{2}$. If $\epsilon_{1}>\epsilon_{2}$, the local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ acting on the hole, i.e., inducing hole spin flips (Eq. (15)), is aligned in positive $z$-direction, while the local electric field acting on the electron, i.e., inducing electron spin flips (Eq. (16)), is aligned in negative $z$-direction. If $\epsilon_{1}<\epsilon_{2}$, the alignments switch. This is a direct consequence of the image charges, which are induced on both dielectrica $\epsilon_{1}$ and $\epsilon_{2}$ surrounding the corresponding real electron or hole charges in the atomically thin semiconductor: If $\epsilon_{1}>\epsilon_{2}$, ($\epsilon_{1}<\epsilon_{2}$) image charges in the dielectricum $\epsilon_{1}$ experience a stronger (weaker) screening compared to image charges in the dielectricum $\epsilon_{2}$, cf. Fig. 4 and Fig. 4. Consequently, an external electric field $E_{z}^{\text{ext}}$ in positive (negative) $z$-direction can add up to (counteracting) the local field acting on the holes $E_{z,\mathbf{q}}^{\text{loc}}$ while counteracting (adding up to) the local electric field on the electrons $E_{z,\mathbf{q}}^{\text{loc}}$, cf. Eq. (17).
Figure 5: Excitonic Rashba matrix element with an external electric field $E_{z}^{\text{ext}}$ (shaded areas) and with a local electric field $E_{z,\mathbf{q}}$ from Eq. (16) (solid lines) for different interlayer distances $h$ in a MoSe2 monolayer on a sapphire substrate.
In Fig. 5, we depict the COM momentum-dependent excitonic Rashba matrix element due to a local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ (solid lines) compared to the Rashba matrix element due to an external electric field $E_{z}^{\text{ext}}$ from Eq. (16) (shaded areas) for different substrate separations $h$, cf. Fig. 1. The external electric field is chosen as such, that the overall magnitude of the Rashba coupling elements of both corresponding fields, external and local, coincide.
It turns out, that the magnitude of the Rashba interaction for a local electric field $E_{z,\mathbf{q}}^{\text{loc}}\neq 0$ at h=0h=0\,nm (black solid line) corresponds to an external electric field of $E_{z}^{\text{ext}}=0.076\,$V nm-1 (grey-shaded area), and, for a realistic substrate distance of h=0.3h=0.3\,nm Florian et al. (2018), the magnitude of the Rashba interaction due to a local electric field $E_{z,\mathbf{q}}^{\text{loc}}\neq 0$ (red solid line) corresponds to an external field of $E_{z}^{\text{ext}}=0.027$ V nm-1 (blue-shaded area). These values, while already substantial, are much lower than the $\mathbf{q}=\mathbf{0}$-values of the local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ at approximately 1 V nm-1 in Fig. 2.
Overall, the magnitude of the excitonic Rashba matrix element for a local electric field decreases by increasing interlayer separation. This behavior stems from the observation from Fig. 2, where a larger interlayer distance leads to a stronger localization of the local electric field in momentum space. This, in turn, reduces the overlap with the excitonic wave functions $\varphi_{\mu,\mathbf{q}}^{K,K,s,s^{\prime}}$, cf. Eq. (16), which exhibit an opposite behavior, i.e., a weaker localization in momentum space, with respect to an increasing interlayer separation due to decreased substrate screening.
IV Equations of Motion for Exciton-Phonon Scattering
To activate the Rashba spin-flip process, generating excitonic occupations $N_{\mu,\mathbf{Q}}$ at $\mathbf{Q}\neq\mathbf{0}$ are necessary, cf. Fig. 5 and Eq. (15) and Eq. (16). For this, we choose exciton-phonon scattering as the most relevant scattering mechanism to generate incoherent excitonic occupations $N_{\mu,\mathbf{Q}}$ Selig et al. (2018); Brem et al. (2020); Thränhardt et al. (2000); Katzer et al. (2023):
$N_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}=\left\langle{{}\hat{P}^{\dagger}}{\vphantom{P}}_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}{{}\hat{P}^{\vphantom{\dagger}}}{\vphantom{P}}_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}\right\rangle_{\text{c}},$ (18)
at above-cryogenic temperatures after optical excitation. To formulate the theory in an appropriate basis for exciton-phonon scattering, we perform a diagonalization of the excitonic Rashba Hamiltonian from Eq. (LABEL:eq:RashbaHamiltonianIntrinsicExciton) into the spin-diagonal basis $s,s^{\prime}\rightarrow\mathcal{S}$ analog to Ref. Mittenzwey et al. (2025). Here, $\mathcal{S}$ is the excitonic spin-diagonal quantum number, which emerges by diagonalizing the excitonic spin eigenvalue equation. For details, we refer to App. C. By adding the exciton-phonon Hamiltonian, the Boltzmann-type equations Katzer et al. (2023); Selig et al. (2018) in the spin-hybrid basis for $\mu=1s$ excitonic states can be derived Mittenzwey et al. (2025):
$\begin{split}\partial_{t}N_{\mathbf{Q}}^{\xi,\xi^{\prime},\mathcal{S}}=&\,\sum_{\begin{subarray}{c}\mathbf{K},\mathcal{S}^{\prime},\\ \xi^{\prime\prime},\xi^{\prime\prime\prime}\end{subarray}}\left(\Gamma^{\text{in},\xi^{\prime\prime},\xi^{\prime\prime\prime},\mathcal{S}^{\prime},\xi,\xi^{\prime},\mathcal{S}}_{\mathbf{K},\mathbf{Q}}N_{\mathbf{K}}^{\xi^{\prime\prime},\xi^{\prime\prime\prime},\mathcal{S}^{\prime}}\right.\\ &\,\quad\quad\quad-\left.\Gamma^{\text{out},\xi^{\prime\prime},\xi^{\prime\prime\prime},\mathcal{S}^{\prime},\xi,\xi^{\prime},\mathcal{S}}_{\mathbf{K},\mathbf{Q}}N_{\mathbf{Q}}^{\xi,\xi^{\prime},\mathcal{S}}\right).\end{split}$ (19)
Here, $\Gamma^{\text{in}/\text{out},j,i}_{\mathbf{K},\mathbf{Q}}$ denote the phonon-assisted scattering rates between states $\mathbf{Q},i$ and $\mathbf{K},j$, which encode intra- ($\xi^{\prime\prime}=\xi$ and $\xi^{\prime\prime\prime}=\xi^{\prime}$) as well as intervalley ($\xi^{\prime\prime}\neq\xi$ or $\xi^{\prime\prime\prime}\neq\xi^{\prime}$) scattering of the corresponding holes and electrons, respectively, forming the exciton. All phonon-related parameters such as phonon modes and electron-phonon coupling elements in deformation potential approximation for the corresponding materials are taken from DFT calculations Li et al. (2013); Jin et al. (2014). Due to Rashba coupling-induced mixing of electron/hole spins in the conduction/valence bands, reflected by the excitonic spin-diagonal quantum number $\mathcal{S}$, phonon-assisted scattering in Eq. (19) does not conserve the spin anymore and phonon-assisted spin relaxation can occur. We note, that the excitonic spin-diagonal quantum number $\mathcal{S}$ corresponds to the old spins $s,s^{\prime}$ at zero center-of-mass momentum $\mathbf{Q}\rightarrow\mathbf{0}$ or in case of a vanishing out-of-plane electric field $E_{z,\mathbf{q}}^{\text{tot}}\rightarrow 0$: $\mathcal{S}\rightarrow s,s^{\prime}$.
From the spin-hybrid excitonic occupations $N_{\mathbf{Q}}^{\xi,\xi^{\prime},\mathcal{S}}$ obtained by solving Eq. (19), the total momentum-integrated excitonic occupations in the old spin basis $N^{\xi,\xi^{\prime},s,s^{\prime}}$ can be retrieved:
$N^{\xi,\xi^{\prime},s,s^{\prime}}=\frac{1}{\mathcal{A}}\sum_{\mathbf{Q},\mathcal{S}}\left|\mathcal{C}_{\mathbf{Q},\mathcal{S}}^{\xi,\xi^{\prime},s,s^{\prime}}\right|^{2}N_{\mathbf{Q}}^{\xi,\xi^{\prime},\mathcal{S}}.$ (20)
Here, $\mathcal{C}_{\mathbf{Q},\mathcal{S}}^{\xi,\xi^{\prime},s,s^{\prime}}$ are the excitonic spin eigenfunctions solving Eq. (44), cf. App. C.
Since the focus of this work is the isolated effect of single phonon-assisted spin flip relaxation due to external or dielectric-environment-induced fields, we neglect double spin flip-inducing intervalley exchange Selig et al. (2020, 2019) and other Coulomb-mediated intervalley scattering mechanisms such as Dexter coupling Berghäuser et al. (2018) linear in the exciton density $N$. However, our theory can be straightforwardly extended to account for these mechanisms as well, similar to the exciton-phonon interaction described above. Also, all density-dependent mechanisms such as Auger scattering Erkensten et al. (2021); Steinhoff et al. (2021) are neglected, since we stay in linear order of the exciton density $N$.
In the following, we examine the local-electric-field-induced phonon-assisted spin relaxation in one example atomically thin semiconductor, a monolayer MoSe2.
V Spin Relaxation in Monolayer MoSe2
We start the analysis with a monolayer MoSe2, where the momentum-direct spin-dark exciton is energetically located at approximately 1.451.45\,meV Lu et al. (2019); Robert et al. (2020) above the optically addressable spin-bright exciton. This small bright-dark splitting is a direct consequence of excitonic effects, as the bare conduction band splitting is on the order of 20 meV Kormányos et al. (2015). Within the used effective-mass approach, this implies a short-range intravalley exchange blue shift Qiu et al. (2015); Yu and Wu (2014) of the spin-bright momentum-direct, i.e., intravalley $K$-$K$ A1s exciton of 5 meV, from which we can estimate the short-range intravalley exchange blue shift of the spin-bright momentum-indirect, i.e., intervalley $KK^{\prime}$ exciton of 9 meV Li et al. (2022); Yang et al. (2022); He et al. (2020). In Fig. 6, we depict the energies at zero center-of-mass momentum of the four considered exciton species.
Due to the small bright-dark splitting of the $K$-$K$ excitons allowing efficient spin hybridization, a monolayer MoSe2 is a promising candidate for efficient excitonic Rashba spin relaxation induced by out-of-plane electric fields. In our case, the electric field is a local field $\mathbf{E}_{z,\mathbf{q}}^{\text{loc}}$ self-consistently induced by the spatial asymmetries of the dielectric environment of the MoSe2 monolayer sandwiched between vacuum and a SiO2 substrate, cf. Eq. (10).
In Fig. 7, we depict the full momentum-dependent exciton dispersions of the lowest spin-bright and spin-dark excitons at the $K$ valley, where the green color gradient denotes the degree of the spin hybridization of the electronic part of the excitons. Due to the slightly smaller exciton mass of the spin-bright exciton and the small bright-dark splitting, a hybridization ”hot spot” around a COM momentum of 0.8 nm-1 occurs. At this region, the initially non-hybridized bands cross, which translates into an anti-crossing, i.e., level repulsion behavior, as soon as a spin-hybridizing electric field is present. The grey solid lines denote Boltzmann distributions at various temperatures, which are important for the interpretation of the exciton dynamics later on.
Figure 6: Energies at zero COM $\mathbf{Q}=\mathbf{0}$ for the four considered exciton configurations.
Figure 7: Dispersions of the spin-bright $E_{\mathbf{Q}}^{K,K,\uparrow,\uparrow}$ (red solid line) and spin-dark $E_{\mathbf{Q}}^{K,K,\uparrow,\downarrow}$ (blue solid line) excitons at the $K$ valley in a MoSe2 monolayer on a SiO2 substrate, cf. also Fig. 9. The green color gradient denotes the degree of the electron spin hybridization via a local electric field. Grey solid lines denote Boltzmann distributions of the spin-like excitonic occupation at various temperatures.
In the numerical simulations, we take into account spin-bright and spin-dark intra- ($N_{\mathbf{Q}}^{K,K,\uparrow,\uparrow}$, $N_{\mathbf{Q}}^{K,K,\uparrow,\downarrow}$) and intervalley ($N_{\mathbf{Q}}^{K,K^{\prime},\uparrow,\uparrow}$, $N_{\mathbf{Q}}^{K,K^{\prime},\uparrow,\downarrow}$) excitonic occupations, cf. also Fig. 6 and Fig. 9, in a spin-hybrid description, cf. Eq. (19). We neglect any hole spin flips, as they are suppressed due to the large A-B splitting. We note, that Rashba coupling between distinct valleys $K$ and $K^{\prime}$ (intervalley spin hybridization) is also strongly suppressed, cf. the discussion to Fig. 2, and therefore neglected. However, intervalley phonon-assisted spin flips are still possible due to the softening of the spin selection rules of the exciton-phonon scattering via intravalley spin hybridization.
As the initial condition for our numerical simulations, we assume a normalized spin-bright excitonic occupation at the $K$ valley within the light cone:
$N_{\mathbf{Q}}^{K,K,\uparrow,\uparrow}(t=0)=\delta_{\mathbf{Q},\mathbf{0}},$ (21)
which roughly approximates the situation after optically exciting the A1s excitonic transition at the $K$ valley with $\sigma_{+}$-light.
Exciton density
Figure 8: Dynamics of the total excitonic occupations $N^{\xi,\xi^{\prime},s,s^{\prime}}$ from Eq. (20) for a MoSe2 monolayer on a SiO2 substrate with $\epsilon_{1}=3.9$ and $\epsilon_{2}=1$: (a) At a temperature of 77 K and a local field manually set as zero, (b) at 77 K, (c) at 150 K and (d) at 300 K.
In Fig. 8, we display the momentum-integrated excitonic occupations $N^{\xi,\xi^{\prime},s,s^{\prime}}$ from Eq. (20), which are obtained by projecting the spin-hybrid solutions of Eq. (19) into the old spin basis, for a MoSe2 monolayer on a SiO2 substrate for a zero local field $E_{z,\mathbf{q}}^{\text{loc}}=0$ at T=77T=77\,K, Fig. 8(a), and for a nonzero local field $E_{z,\mathbf{q}}^{\text{loc}}\neq 0$ at increasing temperatures, Fig. 8(b–d). We do not apply any external fields, i.e., $E_{z}^{\text{ext}}=0$ always holds. At initial time, the optically addressable, momentum-direct spin-bright excitonic occupation is populated (red solid line), which reflects our initial conditions in Eq. (21).
Figure 9: Phonon-assisted spin-conserving intervalley scattering (a), spin-nonconserving intravalley scattering (b) and spin-nonconserving intervalley scattering (c) in a MoSe2 monolayer.
First, we consider the zero-electric-field case ($E_{z,\mathbf{q}}^{\text{loc}}=0$, Eq. (17)) depicted in Fig. 8(a). Here, the spin-bright excitonic intravalley occupation $N^{K,K,\uparrow,\uparrow}$ scatters into the spin-dark intervalley occupation $N^{K,K^{\prime},\uparrow,\uparrow}$ (red dotted line) via spin-conserving phonon-assisted intervalley scattering $N^{K,K,\uparrow,\uparrow}\rightarrow N^{K,K^{\prime},\uparrow,\uparrow}$, cf. process (1) in Fig. 9(a) within 6 ps at 77 K. Spin-dark excitonic occupations remain zero.
Next, in Fig. 8(b), we allow for a nonzero local electric field $E_{z,\mathbf{q}}^{\text{loc}}\neq 0$ in Eq. (17), which emerges due to the asymmetric dielectric structure. This way, next to spin-conserving intervalley scattering, phonon-assisted single spin-nonconserving intravalley scattering into $N^{K,K,\uparrow,\downarrow}$ (blue solid line) within 6 ps is activated, cf. process (3) in Fig. 9(b). Additionally, spin-dark intervalley occupations $N^{K,K^{\prime},\uparrow,\downarrow}$ (blue dotted line) emerge via two paths: On the one hand, it involves two two-step processes. The first two-step process occurs via intervalley spin-conserving scattering into $N^{K,K^{\prime},\uparrow,\uparrow}$, cf. process (1) in Fig. 9(a), and a subsequent intravalley spin-nonconserving scattering into $N^{K,K^{\prime},\uparrow,\downarrow}$, cf. process (4) in Fig. 9(b). The second occurring two-step process involves intravalley spin-nonconserving scattering first, cf. process (3) in Fig. 9(b), and intervalley spin-conserving scattering afterwards, cf. process (2) in Fig. 9(a). On the other hand, spin-nonconserving intervalley scattering from $N^{K,K,\uparrow,\uparrow}$ into $N^{K,K^{\prime},\uparrow,\downarrow}$ also occurs as a one-step process due to the combined action of valley-momentum phonons and spin hybridization, cf. process (5) in Fig. 9(c). However, even though both scattering paths contribute, the first path is significantly more efficient.
Overall, the spin-dark excitonic occupations $N^{K,K,\uparrow,\downarrow}$ and $N^{K,K^{\prime},\uparrow,\downarrow}$ rise on a similar timescale as the spin-conserving intervalley scattering, cf. Fig. 8(a), until an overall equilibrium after 8 ps is reached. The reason for this comparably fast equilibration is the small excitonic bright-dark splitting of 1.45 meV, which renders spin hybridization and therefore spin relaxation very efficient.
At higher temperatures such as 150 K in Fig. 8(c) and 300 K in Fig. 8(d), the overall spin-conserving intervalley and spin-nonconserving intra- as well as intervalley scattering dramatically increases: At 150 K an overall equilibrium is reached after 1 ps, which speeds up to 0.2 ps at 300 K. This overall speed up occurs due to thermal activation of phonons and a thermalization of excitonic occupations. The former increases the induced absorption and emission from and into the phonon bath at increasing temperature and the latter results in a temperature-dependent broadening of the (Boltzmann) distribution in momentum space $\mathbf{Q}$, cf. grey solid lines in Fig. 7. The latter phenomenon is especially responsible for a speed-up of the spin relaxation, as excitonic states at larger center-of-mass momenta $\mathbf{Q}$ become increasingly populated, where the spin relaxation is more efficient due to the momentum-dependent spin coupling, cf. Fig. 5, and, most importantly, due to an increasing probability of scattering via the ”hot spot”, cf. Fig. 7. In Tab. 1, we depict the explicit spin relaxation times, which are obtained by extracting the risetime of the spin-dark intravalley occupations at the $K$ valley from the numerical simulations in Fig. 8.
77 K 150 K 300 K
MoSe2 2.2 ps 0.32 ps 88 fs
MoS2 5 ns 0.8 ns 25 ps
Table 1: Spin relaxation times for spin flips of intravalley excitons for MoSe2 and MoS2 monolayers on a SiO2 substrate in units of ps-1. The spin relaxation times are obtained by extracting the risetime of spin-dark intravalley occupations at the $K$ valley from numerical simulations.
However, care should be taken in comparing specific numbers to distinct experiments, as a specific dielectric environment has been chosen for the calculation. Also, the bright-dark splitting plays a critical role. To illustrate this, we have also performed numerical simulations for a MoS2 monolayer for the same dielectric environment. The extracted spin relaxation times are displayed in Tab. 1 and are 5 ns at 77 K, which speed up to 25 ps at 300 K. Compared to a MoSe2 monolayer, they are orders of magnitude slower, which is a direct consequence of the larger bright-dark splitting of 14 meV Robert et al. (2020) compared to 1.45 meV in MoSe2 rendering the spin hybridization much less efficient. Other monolayer materials such as WSe2 or WSe2 behave similar, as their bright-dark splitting is even larger (40–55 meV Molas et al. (2019); Mapara et al. (2022); Molas et al. (2017); Zhou et al. (2017)). In these materials, other spin relaxation mechanisms such as spin-nonconserving scattering via chiral phonons Lagarde et al. (2024); Chan et al. (2025); Zhang and Cao (2022) are expected to dominate over a Rashba spin-orbit interaction induced by spatial asymmetries of the dielectric environment. However, in TMDC heterostructures such as MoSe2/MoS2 bilayers, intra-exciton fields due to charge-separated interlayer exciton densities can induce a significant many-body Rashba effect Mittenzwey et al. (2025). Thus, the exact choice of the material and dielectric environment conditions determine, which spin relaxation mechanism is dominant.
VI Conclusion
In this work, we discovered a Rashba single spin-flip mechanism due to local out-of-plane electric fields induced by spatial asymmetries in the dielectric environment. We developed an excitonic many-body description, explicitly calculated the local electric fields for a given dielectric environment and quantified the Rashba coupling strength. For a monolayer MoSe2, we demonstrated very efficient spin relaxation on a sub-picosecond (150 K) to a hundred-femtosecond (300 K) timescale via a phonon-assisted single electron spin-flip channel.
Acknowledgements.
H.M. and A.K. acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG) through Project KN 427/11-2 Project No. 420760124. A.K. acknowledges financial support by the DFG through Project KN 427/15-1, Project No. 556436549. We thank M. Snoeken for his help. H.M. thanks M. Selig for inspiration.
Appendix A Derivation of the Many-Particle Rashba Hamiltonian
Classical spin-orbit Hamiltonian Winkler (2003):
$H_{\text{BR}}^{(1)}=-\frac{\alpha_{\text{BR}}}{\hbar}\bm{\sigma}\cdot\left(\mathbf{E}(\mathbf{r},z)\times\mathbf{p}\right).$ (22)
The second-quantized Hamiltonian is obtained via an expansion in Bloch wave functions:
$\psi_{\mathbf{k},\lambda}^{s}(\mathbf{r},z)=\frac{1}{\sqrt{\mathcal{A}}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}u_{\lambda,\mathbf{k}}(\mathbf{r})\xi(z)\chi_{s},$ (23)
and reads:
$\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2},\\ s_{1},s_{2}\end{subarray}}S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\mathbf{k}_{1},\lambda_{1}}^{s_{1}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}},$ (24)
where the Rashba matrix element is given by:
$S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}=\int\mathrm{d}^{3}r\,{\psi}^{*}{\vphantom{\psi}}_{\mathbf{k}_{1},\lambda_{1}}^{s_{1}}(\mathbf{r})H_{\text{BR}}^{(1)}\psi{\vphantom{\psi}}_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}}(\mathbf{r}).$ (25)
With the wave functions from Eq. (23), it follows:
$\begin{split}S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}=&\,-\frac{\alpha_{\text{BR}}}{\hbar}\frac{1}{\mathcal{A}}\int\mathrm{d}^{2}r\,\mathrm{d}z\,\mathrm{e}^{-\mathrm{i}\mathbf{k}_{1}\cdot\mathbf{r}}u^{*}\vphantom{u}_{\mathbf{k}_{1},\lambda_{1}}{}^{s_{1}}(\mathbf{r})\xi^{*}(z)\chi_{s_{1}}^{*}\hat{E}_{z}(\mathbf{r},z)\left(\sigma_{y}p_{x}-\sigma_{x}p_{y}\right)\mathrm{e}^{\mathrm{i}\mathbf{k}_{2}\cdot\mathbf{r}}u_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}}(\mathbf{r})\xi(z)\chi_{s_{2}},\end{split}$ (26)
with the Rashba coupling constant $\alpha_{\text{BR}}$ in units of $\alpha_{\text{BR}}]=e\,\mathrm{nm}^{2}$. Writing the electric field in Fourier representation $\hat{E}_{z}(\mathbf{r},z)=\frac{1}{\mathcal{A}}\sum_{\mathbf{q},\mathbf{G}}\mathrm{e}^{\mathrm{i}(\mathbf{q}+\mathbf{G})\cdot\mathbf{r}}\hat{E}_{z,\mathbf{q}+\mathbf{G}}(z)$ Wiser (1963), where $\mathbf{q}$ is a crystal momentum within the first Brillouin zone and $\mathbf{G}$ is a reciprocal lattice vector, and identifying the quantum-confined electric field as:
$\hat{E}_{z,\mathbf{q}}=\frac{1}{\mathcal{A}}\int\,\mathrm{d}z\,\hat{E}_{z,\mathbf{q}}(z)|\xi(z)|^{2},$ (27)
yields:
$\begin{split}S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}=-\frac{\alpha_{\text{BR}}}{\hbar}\sum_{\mathbf{q},\mathbf{G}}\frac{1}{\mathcal{A}}\hat{E}_{z,\mathbf{q}+\mathbf{G}}\int\mathrm{d}^{2}r\,\mathrm{e}^{\mathrm{i}(\mathbf{q}+\mathbf{G})\cdot\mathbf{r}}\mathrm{e}^{-\mathrm{i}\mathbf{k}_{1}\cdot\mathbf{r}}u^{*}\vphantom{u}_{\mathbf{k}_{1},\lambda_{1}}{}^{s_{1}}(\mathbf{r})\chi_{s_{1}}^{*}\left(\sigma_{y}p_{x}-\sigma_{x}p_{y}\right)\chi_{s_{2}}\mathrm{e}^{\mathrm{i}\mathbf{k}_{2}\cdot\mathbf{r}}u_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}}(\mathbf{r}).\end{split}$ (28)
Letting the momentum operator act on the Bloch wave function yields:
$\begin{split}S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}=-\frac{\alpha_{\text{BR}}}{\hbar}\sum_{\mathbf{q},\mathbf{G}}\frac{1}{\mathcal{A}}\hat{E}_{z,\mathbf{q}+\mathbf{G}}\int\mathrm{d}^{2}r\,\mathrm{e}^{-\mathrm{i}(\mathbf{k}_{1}-\mathbf{k}_{2}-\mathbf{q}+\mathbf{G})\cdot\mathbf{r}}u^{*}\vphantom{u}_{\mathbf{k}_{1},\lambda_{1}}{}^{s_{1}}(\mathbf{r})\chi_{s_{1}}^{*}\left(\sigma_{y}(\hbar k_{2,x}+p_{x})-\sigma_{x}(\hbar k_{2,y}+p_{y})\right)\chi_{s_{2}}u_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}}(\mathbf{r}).\end{split}$ (29)
Expanding the integral over the unit cells yields:
$S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}=-\frac{\alpha_{\text{BR}}}{\hbar}\sum_{\mathbf{q},\mathbf{G}}\frac{1}{\mathcal{A}}\hat{E}_{z,\mathbf{q}+\mathbf{G}}\sum_{\mathbf{R}}\mathrm{e}^{-\mathrm{i}(\mathbf{k}_{1}-\mathbf{k}_{2}-\mathbf{q}-\mathbf{G})\cdot\mathbf{R}}\\ \times\int_{\mathcal{A}_{\text{UC}}}\mathrm{d}^{2}r\,\mathrm{e}^{-\mathrm{i}(\mathbf{k}_{1}-\mathbf{k}_{2}-\mathbf{q}-\mathbf{G})\cdot\mathbf{r}}u^{*}\vphantom{u}_{\mathbf{k}_{1},\lambda_{1}}{}^{s_{1}}(\mathbf{r})\chi_{s_{1}}^{*}\left(\sigma_{y}(\hbar k_{2,x}+p_{x})-\sigma_{x}(\hbar k_{2,y}+p_{y})\right)\chi_{s_{2}}u_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}}(\mathbf{r}).$ (30)
The lattice sum yields:
$\sum_{\mathbf{R}}\mathrm{e}^{-\mathrm{i}(\mathbf{k}_{1}-\mathbf{k}_{2}-\mathbf{q}-\mathbf{G})\cdot\mathbf{R}}=N\delta_{\mathbf{k}_{1},\mathbf{k}_{2}+\mathbf{q}+\mathbf{G}},$ (31)
where $N$ is the total number of unit cells. Considering further the action of the Pauli matrices on the spin eigenfunctions:
$\sigma_{x}\chi_{s}=\chi_{\bar{s}},\quad\sigma_{y}\chi_{s}=\mathrm{i}\left(\delta_{s,\uparrow}-\delta_{s,\downarrow}\right)\chi_{\bar{s}},$ (32)
yields:
$S_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}=-\frac{\alpha_{\text{BR}}}{\hbar}\sum_{\mathbf{q},\mathbf{G}}\hat{E}_{z,\mathbf{q}+\mathbf{G}}\delta_{\mathbf{k}_{1},\mathbf{k}_{2}+\mathbf{q}+\mathbf{G}}\delta_{s_{1},\bar{s}_{2}}\\ \times\frac{1}{\mathcal{A}_{\text{UC}}}\int_{\mathcal{A}_{\text{UC}}}\mathrm{d}^{2}r\,\mathrm{e}^{\mathrm{i}\mathbf{G}\cdot\mathbf{r}}u^{*}\vphantom{u}_{\mathbf{k}_{2}+\mathbf{q},\lambda_{1}}{}^{\bar{s}_{2}}(\mathbf{r})\left(\mathrm{i}\left(\delta_{s,\uparrow}-\delta_{s,\downarrow}\right)(\hbar k_{2,x}+p_{x})-(\hbar k_{2,y}+p_{y})\right)u_{\mathbf{k}_{2},\lambda_{2}}^{s_{2}}(\mathbf{r}).$ (33)
Identifying the momentum matrix element as:
$\mathbf{p}_{\mathbf{k}_{1},\mathbf{k}_{2},\lambda_{1},\lambda_{2}}^{s_{1},s_{2}}(\mathbf{G})=\frac{1}{\mathcal{A}_{\text{UC}}}\int_{\mathcal{A}_{\text{UC}}}\,u^{*}\vphantom{u}_{\lambda_{1},\mathbf{k}_{1}}{}^{s_{1}}(\mathbf{r})\mathrm{e}^{\mathrm{i}\mathbf{G}\cdot\mathbf{r}}\left(\hbar\mathbf{k}_{2}+\mathbf{p}\right)u_{\lambda_{2},\mathbf{k}_{2}}^{s_{2}}(\mathbf{r}),$ (34)
expanding over the band extrema with $\mathbf{k}\rightarrow\mathbf{k}+\mathbf{K}^{\xi}$ and $\mathbf{q}\rightarrow\mathbf{q}+\mathbf{K}^{\xi^{\prime}}$ and neglecting Umklapp processes with $\mathbf{G}\neq\mathbf{0}$ yields Eq. (2) in the main manuscript.
Appendix B Calculation of the Local Electric Field
Starting point is the generalized Poisson equation for the Coulomb potential $\hat{\phi}$:
$\nabla\cdot\left(\epsilon(\mathbf{r})\nabla\phi(\mathbf{r})\right)=-\hat{\rho}(\mathbf{r}).$ (35)
Redefining $\mathbf{r}\equiv\left(\mathbf{r},z\right)$ and applying a Fourier transformation for the in-plane coordinates leads to:
$-|\mathbf{q}|^{2}\hat{\phi}_{\mathbf{q}}(z)+\frac{1}{\epsilon_{\parallel}(z)}\partial_{z}\left(\epsilon_{\bot}(z)\partial_{z}\hat{\phi}_{\mathbf{q}}(z)\right)=-\frac{1}{\epsilon_{\parallel}(z)}\hat{\rho}_{\mathbf{q}}(z).$ (36)
Here, the inhomogeneity is caused by the operator-valued charge density from Eq. (7). We reformulate the Poisson equation for the Green’s function $G_{\mathbf{q}}(z,z^{\prime})$, defined by $\hat{\phi}_{\mathbf{q}}(z)=\int_{-\infty}^{\infty}\mathrm{d}z^{\prime}\,G_{\mathbf{q}}(z,z^{\prime})\hat{\rho}_{\mathbf{q}}(z^{\prime})$:
$-|\mathbf{q}|^{2}G_{\mathbf{q}}(z,z^{\prime})+\frac{1}{\epsilon_{\parallel}(z)}\partial_{z}\left(\epsilon_{\bot}(z)\partial_{z}G_{\mathbf{q}}(z,z^{\prime})\right)=-\frac{1}{\epsilon_{\parallel}(z)}\delta(z-z^{\prime}).$ (37)
By defining the Poisson equation piecewise in a geometry of five dielectric slabs (we assume a small interlayer gap $h$ between substrate and superstrate Florian et al. (2018)), where the inhomogeneity is located in the middle layer, and employing the electrostatic boundary conditions, the Green’s function can be solved for:
$\begin{split}&G_{\mathbf{q}}(z,z^{\prime})=\frac{1}{2\epsilon_{0}\epsilon_{s}|\mathbf{q}|}\mathrm{e}^{-\bar{\epsilon}_{s}|\mathbf{q}||z-z^{\prime}|}\\ &\,+\frac{1}{2\epsilon_{0}\epsilon_{s}|\mathbf{q}|f_{\mathbf{q}}}\left(\left(\epsilon_{s,-}\epsilon_{2,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{s,+}\epsilon_{2,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\left(\epsilon_{1,+}\epsilon_{s,-}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{s,+}\mathrm{e}^{-h|\mathbf{q}|}\right)\mathrm{e}^{-d|\mathbf{q}|}\left(\mathrm{e}^{|\mathbf{q}|(z-z^{\prime})}+\mathrm{e}^{-|\mathbf{q}|(z-z^{\prime})}\right)\right.\\ &\,\left.+\left(\epsilon_{s,-}\epsilon_{2,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{s,+}\epsilon_{2,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\left(\epsilon_{1,+}\epsilon_{s,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{s,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\mathrm{e}^{|\mathbf{q}|(z+z^{\prime})}\right.\\ &\,\left.+\left(\epsilon_{s,+}\epsilon_{2,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{s,-}\epsilon_{2,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\left(\epsilon_{1,+}\epsilon_{s,-}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{s,+}\mathrm{e}^{-h|\mathbf{q}|}\right)\mathrm{e}^{-|\mathbf{q}|(z+z^{\prime})}\right).\end{split}$ (38)
The electric field, which is generated by the Coulomb potential $\phi$, is then obtained by:
$\hat{E}_{z,\mathbf{q}}(z)=-\partial_{z}\hat{\phi}_{\mathbf{q}}(z)=-\partial_{z}\int\mathrm{d}z^{\prime}\,G_{\mathbf{q}}(z,z^{\prime})\hat{\rho}_{\mathbf{q}}(z^{\prime}).$ (39)
By performing the $z$-derivative, separating the $z$-dependence of the charge density $\hat{\rho}_{\mathbf{q}}(z)$ and employing confinement wave functions of the form:
$\xi(z)=\sqrt{\frac{2}{d}}\cos\left(\frac{\pi}{d}z\right),\quad-\frac{d}{2}\leq z\leq\frac{d}{2},$ (40)
the $z$- and $z^{\prime}$-integrals in Eq. (LABEL:eq:ElectricFieldConfined) can be performed, which yields Eq. (10), where:
$\begin{split}g_{\mathbf{q}}=&\,\left(\left(\epsilon_{s,-}\epsilon_{2,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{s,+}\epsilon_{2,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\left(\epsilon_{1,+}\epsilon_{s,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{s,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\right.\\ &\,\left.-\left(\epsilon_{s,+}\epsilon_{2,+}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{s,-}\epsilon_{2,-}\mathrm{e}^{-h|\mathbf{q}|}\right)\left(\epsilon_{1,+}\epsilon_{s,-}\mathrm{e}^{h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{s,+}\mathrm{e}^{-h|\mathbf{q}|}\right)\right)\left(\frac{8\pi^{2}\sinh\left(\frac{\bar{\epsilon}_{s}|\mathbf{q}|d}{2}\right)}{\bar{\epsilon}_{s}|\mathbf{q}|d\left(4\pi^{2}+\left(\bar{\epsilon}_{s}|\mathbf{q}|d\right)^{2}\right)}\right)^{2},\end{split}$ (41)
and:
$\begin{split}f_{\mathbf{q}}=&\,-\epsilon_{s,-}\epsilon_{s,-}\left(\epsilon_{1,+}\epsilon_{2,+}\mathrm{e}^{-\bar{\epsilon}_{s}d|\mathbf{q}|}\mathrm{e}^{2h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{2,-}\mathrm{e}^{\bar{\epsilon}_{s}d|\mathbf{q}|}\mathrm{e}^{-2h|\mathbf{q}|}\right)\\ &\,\left.-2\epsilon_{s,+}\epsilon_{s,-}\left(\epsilon_{1}\epsilon_{2}-1\right)\left(\mathrm{e}^{\bar{\epsilon}_{s}d|\mathbf{q}|}-\mathrm{e}^{-\bar{\epsilon}_{s}d|\mathbf{q}|}\right)\right.\\ &\,+\epsilon_{s,+}\epsilon_{s,+}\left(\epsilon_{1,+}\epsilon_{2,+}\mathrm{e}^{\bar{\epsilon}_{s}d|\mathbf{q}|}\mathrm{e}^{2h|\mathbf{q}|}-\epsilon_{1,-}\epsilon_{2,-}\mathrm{e}^{-\bar{\epsilon}_{s}d|\mathbf{q}|}\mathrm{e}^{-2h|\mathbf{q}|}\right),\end{split}$ (42)
with:
$\epsilon_{i,\pm}=\epsilon_{i}\pm 1,\quad\epsilon_{i}=\sqrt{\epsilon_{\parallel,i}\epsilon_{\perp,i}},\quad\bar{\epsilon}_{i}=\sqrt{\frac{\epsilon_{\parallel,i}}{\epsilon_{\perp,i}}}.$ (43)
Appendix C Excitonic Spin Eigenvalue Equation
The excitonic spin eigenvalue equation reads Mittenzwey et al. (2025):
$\begin{split}E_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s_{1},s_{2}}\mathcal{C}_{\mu,\mathbf{Q},\mathcal{S}}^{\xi,\xi^{\prime},s_{1},s_{2}}+\sum_{s_{3},s_{4}}\left(S_{\mu,\mathbf{Q}}^{e,\xi,\xi^{\prime},s_{1},s_{2}}\delta^{s_{3},s_{1}}_{s_{4},\bar{s}_{2}}-S_{\mu,\mathbf{Q}}^{h,\xi,\xi^{\prime},s_{1},s_{2}}\delta^{s_{3},\bar{s}_{1}}_{s_{4},s_{2}}\right)\mathcal{C}_{\mu,\mathbf{Q},\mathcal{S}}^{\xi,\xi^{\prime},s_{3},s_{4}}=E_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},\mathcal{S}}\mathcal{C}_{\mu,\mathbf{Q},\mathcal{S}}^{\xi,\xi^{\prime},s_{1},s_{2}},\end{split}$ (44)
[…]
The excitonic transitions are transformed as follows:
${{}\hat{P}^{\vphantom{\dagger}}}{\vphantom{P}}_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},s,s^{\prime}}=\sum_{\mathcal{S}}\mathcal{C}_{\mu,\mathbf{Q},\mathcal{S}}^{\xi,\xi^{\prime},s,s^{\prime}}{{}\hat{P}^{\vphantom{\dagger}}}{\vphantom{P}}_{\mu,\mathbf{Q}}^{\xi,\xi^{\prime},\mathcal{S}},$ (45)
with:
[…]
which enables a formulation of the excitonic theory in the spin-diagonal basis $\mathcal{S}$.
Appendix D Parameters
Quantity MoSe2 MoS2
Bright-dark splitting $E_{1s,\mathbf{Q}=\mathbf{0}}^{K,K,\uparrow,\uparrow}-E_{1s,\mathbf{Q}=\mathbf{0}}^{K,K,\uparrow,\downarrow}$ (meV) 1.45-1.45 Robert et al. (2020); Lu et al. (2019) 14 Robert et al. (2020)
Effective spin-up electron mass $m_{e}^{K,\uparrow}/m_{0}$ Kormányos et al. (2015) 0.495 0.435
Effective spin-down electron mass $m_{e}^{K,\uparrow}/m_{0}$ Kormányos et al. (2015) 0.57 0.465
Effective spin-up hole mass $m_{h}^{K,\uparrow}/m_{0}$ Kormányos et al. (2015) 0.595 0.54
Rashba coupling constant for electrons $\alpha_{\text{BR}}^{e,K}$ ($e$ nm2) Yang et al. (2015) 0.011 0.003
Rashba coupling constant for holes $\alpha_{\text{BR}}^{h,K}$ ($e$ nm2) Yang et al. (2015) 0.0422 0.041
Monolayer width $d$ (nm) Kylänpää and Komsa (2015) 0.6527 0.618
Static out-of-plane dielectric constant $\epsilon_{\bot,0}$ Laturia et al. (2018) 7.2 6.2
Static bulk dielectric constant $\epsilon_{\text{bulk},0}=\sqrt{\epsilon_{\text{bulk},\parallel,0}\epsilon_{\text{bulk},\bot,0}}$ Laturia et al. (2018) 12.05 10.47
Plasmon peak energy $\hbar\omega_{\text{pl}}$ (eV) Kumar and Ahluwalia (2012) 22.0 22.5
Thomas-Fermi parameter $\alpha_{\text{TF}}$ (fit to ab-initio results Andersen et al. (2015)) 1.9 1.5
Calculated $1s$ excitonic binding energies
with $h=0.3$ nm Florian et al. (2018) (meV)
Spin-bright, SiO2 substrate 441-441 478-478
Spin-dark, SiO2 substrate 455-455 486-486
Spin-bright, Al2O3 substrate 341-341 367-367
Spin-dark, Al2O3 substrate 355-355 375-375
Table 2: Parameters used in the simulations.
In Tab. 2, we display the parameters used in all numerical simulations.
For the numerical evaluation of the Wannier equation, which provides the excitonic binding energies and wave functions, we use the analytical dielectric function from Ref. Trolle et al. (2017), which is fitted to ab-initio calculations from the Computational Materials Repository Andersen et al. (2015), and the screened Coulomb potential $V_{\mathbf{q}}=\frac{e^{2}}{\mathcal{A}}\int\mathrm{d}z\,\mathrm{d}z^{\prime}\,|\xi(z)|^{2}G_{\mathbf{q}}(z,z^{\prime})|\xi(z^{\prime})|^{2}$ with the Green’s function from Eq. (38).
All relevant parameters for the exciton-phonon interaction such as electron-phonon interaction potentials, phonon dispersions and phonon modes are taken from Refs. Jin et al. (2014); Li et al. (2013). For intravalley scattering, we take into account $A^{\prime}$, TO, LA and TA modes and for intervalley scattering we take into account $A^{\prime}$, LO, TO, LA and TA modes. The interaction potentials in deformation-potential approximation from Refs. Jin et al. (2014); Li et al. (2013) for acoustic phonons are rescaled by a factor of $\frac{1}{\sqrt{2}}$ to properly consider the fact, that piezoelectric coupling is ineffective for excitons Lengers et al. (2020).

### v2

@@
 Abstract
-We propose a previously unexplored spin-orbit interaction mechanism by establishing a mesoscopic many-particle Rashba Hamiltonian.
+We propose a pair spin-orbit interaction (PSOI) mechanism by establishing a mesoscopic many-particle Rashba Hamiltonian.
 In lowest order, this Hamiltonian self-consistently describes exciton spin relaxation in monolayer transition metal dichalcogenides (TMDC) due to local electric fields caused by spatial asymmetries in the dielectric environment.
@@
 Moreover, single spin flips can occur via, e.g., Rashba interaction.
-Rashba spin-orbit interaction can already affect linear optical spectra of a spin-forbidden mixture of $s$-$p$ orbital excitons in monolayers Cao et al.
+Rashba spin-orbit interaction (SOI) can already affect linear optical spectra of a spin-forbidden mixture of $s$-$p$ orbital excitons in monolayers Cao et al.
 (2024, 2025).
@@
 (2025); Zhang and Cao (2022).
-In this work, we derive a new many-body Rashba spin-orbit interaction mechanism Winkler (2003); Bychkov and Rashba (1984); Kormányos et al. (2014):
+Besides the one-body SOI, it has been shown, that many-body-, Breit- or pair spin-orbit interaction (PSOI) due to self-induced Coulomb fields also occurs Breit (1929); Yi and Sheng (2000); Gindikin and Sablikov (2018, 2022); Liu and Principi (2024); Rajagopal and Mochena (1998), which completes the electron-electron interaction picture.
+In particular, such process can induce a superconducting phase in thin films under suitable conditions Gindikin and Kamenev (2025).
+In this work, we derive a many-body Rashba PSOI mechanism Winkler (2003); Bychkov and Rashba (1984); Kormányos et al. (2014); Gindikin and Sablikov (2022); Chang (1993):
 $H_{\text{BR}}=-\frac{\alpha_{\text{BR}}}{\hbar}\bm{\sigma}\cdot\left(\mathbf{E}(\mathbf{r},z)\times\mathbf{p}\right),$ (1)
-which results from self-consistently described out-of-plane electric fields $\mathbf{E}(\mathbf{r},z)$, induced by the dielectric environment.
-Here, $\bm{\sigma}=\begin{pmatrix}\sigma_{x}&\sigma_{y}&\sigma_{z}\end{pmatrix}^{\top}$ is the vector of the Pauli matrices $\sigma_{i}$, $\mathbf{p}=\frac{\hbar}{\mathrm{i}}\nabla$ is the momentum operator and $\alpha_{\text{BR}}$ is the Rashba coupling constant, which, in solids, is greatly enhanced compared to the atomic constant appearing in the Pauli equation Kormányos et al.
+which results from self-consistently described out-of-plane electric fields $\mathbf{E}(\mathbf{r},z)$, induced by the dielectric environment and examine its consequences for excitons in atomically thin semiconductors.
+Here, $\bm{\sigma}=\begin{pmatrix}\sigma_{x}&\sigma_{y}&\sigma_{z}\end{pmatrix}^{\top}$ is the vector of the Pauli matrices $\sigma_{i}$, $\mathbf{p}=\frac{\hbar}{\mathrm{i}}\nabla$ is the momentum operator and $\alpha_{\text{BR}}$ is the Rashba coupling constant, which, in solids, is greatly enhanced compared to the atomic constant in vacuum Kormányos et al.
 (2014); Winkler (2003).
@@
 In combination with exciton-phonon interaction generating non-zero COM-momentum excitons (Sec.
-IV), the described spin-orbit interaction leads to a phonon-assisted exciton spin relaxation.
+IV), the described PSOI leads to a phonon-assisted exciton spin relaxation.
 We illustrate the new mechanism for a monolayer MoSe2 and quantify its impact on the incoherent exciton dynamics via microscopic calculations (Sec.
@@
 (2):
-$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\\ \lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\\ \xi,\xi^{\prime},\xi^{\prime\prime},s,s^{\prime}\end{subarray}}E_{z,\mathbf{q}+\mathbf{K}^{\xi^{\prime}}-\mathbf{K}^{\xi}}^{\text{loc}}F_{\lambda_{2},\lambda_{3},\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},\xi^{\prime\prime},s^{\prime},s^{\prime}}\\ &\,\quad\times S_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda_{1},\lambda_{4},\xi^{\prime},\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{1},\mathbf{k}+\mathbf{q}}^{\xi^{\prime},\bar{s}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{2},\mathbf{k}^{\prime}-\mathbf{q}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{3},\mathbf{k}^{\prime}}^{\xi^{\prime\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{4},\mathbf{k}}^{\xi,s}.\end{split}$ (9)
+$\begin{split}&\hat{H}_{\text{BR}}=\sum_{\begin{subarray}{c}\mathbf{k},\mathbf{k}^{\prime},\mathbf{q},\\ \lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\\ \xi,\xi^{\prime},\xi^{\prime\prime},s,s^{\prime}\end{subarray}}E_{z,\mathbf{q}+\mathbf{K}^{\xi^{\prime}}-\mathbf{K}^{\xi}}^{\text{loc}}F_{\lambda_{2},\lambda_{3},\mathbf{k}^{\prime}-\mathbf{q},\mathbf{k}^{\prime}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},\xi^{\prime\prime},s^{\prime},s^{\prime}}\\ &\,\quad\times S_{\mathbf{k}+\mathbf{q},\mathbf{k}}^{\lambda_{1},\lambda_{4},\xi^{\prime},\xi,\bar{s},s}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{1},\mathbf{k}+\mathbf{q}}^{\xi^{\prime},\bar{s}}{\hat{a}^{\dagger}}{\vphantom{a}}_{\lambda_{2},\mathbf{k}^{\prime}-\mathbf{q}}^{\xi^{\prime\prime}+\xi-\xi^{\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{3},\mathbf{k}^{\prime}}^{\xi^{\prime\prime},s^{\prime}}{\hat{a}^{\vphantom{\dagger}}}{\vphantom{a}}_{\lambda_{4},\mathbf{k}}^{\xi,s}.\end{split}$ (9)
 Eq.
@@
 This way, the excitonic Rashba Hamiltonian in Eq.
-(LABEL:eq:RashbaHamiltonianIntrinsicExciton) combines two physically distinct fields, a local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ due to many-particle interactions and an arbitrary in-plane spatially homogeneous external electric field $E_{z}^{\text{ext}}$, in one consistent description, which is described by a total electric field $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ entering Eq.
+(LABEL:eq:RashbaHamiltonianIntrinsicExciton) combines two physically distinct fields, a local electric field $E_{z,\mathbf{q}}^{\text{loc}}$ due to many-particle interactions inducing PSOI and an arbitrary in-plane spatially homogeneous external electric field $E_{z}^{\text{ext}}$ inducing SOI, in one consistent description, which is described by a total electric field $E_{z,\mathbf{q}}^{h/e,\text{tot}}$ entering Eq.
 (15) and Eq.
@@
 The green color gradient denotes the degree of the electron spin hybridization via a local electric field.
-Grey solid lines denote Boltzmann distributions of the spin-like excitonic occupation at various temperatures.
+Grey solid lines denote Boltzmann distributions of the spin-bright excitonic occupation at various temperatures.
 In the numerical simulations, we take into account spin-bright and spin-dark intra- ($N_{\mathbf{Q}}^{K,K,\uparrow,\uparrow}$, $N_{\mathbf{Q}}^{K,K,\uparrow,\downarrow}$) and intervalley ($N_{\mathbf{Q}}^{K,K^{\prime},\uparrow,\uparrow}$, $N_{\mathbf{Q}}^{K,K^{\prime},\uparrow,\downarrow}$) excitonic occupations, cf.
@@
 (2024); Chan et al.
-(2025); Zhang and Cao (2022) are expected to dominate over a Rashba spin-orbit interaction induced by spatial asymmetries of the dielectric environment.
+(2025); Zhang and Cao (2022) are expected to dominate over a Rashba PSOI induced by spatial asymmetries of the dielectric environment.
 However, in TMDC heterostructures such as MoSe2/MoS2 bilayers, intra-exciton fields due to charge-separated interlayer exciton densities can induce a significant many-body Rashba effect Mittenzwey et al.
@@
 VI Conclusion
-In this work, we discovered a Rashba single spin-flip mechanism due to local out-of-plane electric fields induced by spatial asymmetries in the dielectric environment.
+In this work, we established a Rashba single spin-flip mechanism due to local out-of-plane electric fields self-consistently induced by spatial asymmetries in the dielectric environment in atomically thin semiconductors.
 We developed an excitonic many-body description, explicitly calculated the local electric fields for a given dielectric environment and quantified the Rashba coupling strength.
@@
 556436549.
-We thank M.
-Snoeken for his help.
+We thank Michiel Snoeken for his help.
 H.M.
-thanks M.
-Selig for inspiration.
+thanks Malte Selig for inspiration.
+Data Availability Statement
+The data supporting the findings of this work are available from the authors within reasonable request.
 Appendix A Derivation of the Many-Particle Rashba Hamiltonian